OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = −3x + 12 and passes through the point (−1, 6). y = one thirdx + 7 y = −3x + 3 y = one thirdx + 3 y = −3x + 7

OpenStudy (anonymous):

you would have to know that parallel lines have the same slope..

OpenStudy (anonymous):

2nd one is right answer.

OpenStudy (anonymous):

Equation of a straight line: y = mx + c Where m is the gradient 2 lines are parallel when the gradient of them both is equal so as long as you have y = -3x then you can put any number after it and the new line would be Parallell e.g. y = -3x + 56434657567099123 y = -3x + 1 For it to pass through ( -1 , 6 ) then put the numbers in and see what you get ( but dont change the gradient ) e.g. y = -3x + 3 ( when x = -1 , y = 6 )

OpenStudy (anonymous):

thanks so much! i have more do you mind?

OpenStudy (anonymous):

nop.

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). y = 2x + 6 y = −one halfx + 3 y = −one halfx + 6 y = 2x + 3

OpenStudy (anonymous):

that is the last question

OpenStudy (anonymous):

3rd one is answer.y = −one halfx + 6

OpenStudy (anonymous):

i cant thank you enough

OpenStudy (anonymous):

i amnot sure about last one sorry.

OpenStudy (anonymous):

thats okay :) i got it right but i still got a 60 on the overall assignment :( i have to take it again.